OpenStudy (anonymous):

Write the equation of the line that satisfies the given conditions. Express the final equation in standard form. Contains the point (-2, 4) and is perpendicular to the line 3x − y = 4

OpenStudy (anonymous):

i hope you don't mind that i solve this in slope intercept form instead.

OpenStudy (anonymous):

I don't mind. as long as I can properly convert it to standard

OpenStudy (anonymous):

3x-y=4, rearrange to get y=3x-4, the slope for this equation is 3. since the wanted equation is perpendicular to this eqation, its slope will be -1/3. plug in (-2,4) in y=mx+b 4=-(1/3)(-2)+b and solve for b

OpenStudy (anonymous):

You put it in standard form?

OpenStudy (anonymous):

yes, after you've figured out what the slope and the intercept are

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 4= 2/3b which reduced is b=2/3

OpenStudy (anonymous):

4=(2/3)+b isolate b b is not equal to 2/3

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

b= 10/3 ?

OpenStudy (anonymous):

no 4=(2/3)+b what do you need to do to isolate b?

OpenStudy (anonymous):

subtract the 2/3

OpenStudy (anonymous):

so -2/3 + 4 = b

OpenStudy (anonymous):

sorry, you were right it is 10/3, I wasn't thinking when i saw that you had written 4= 2/3b where b=2/3

OpenStudy (anonymous):

Its ok.. So with 10/3, where do I go now to solve the problem?

OpenStudy (anonymous):

plug that the slope into the standard form equation